OpenStudy (anonymous):

Shakespeare is buiding a shed with a square base measuring x feet by x feet, and which is h feet high. He wants the volume of the shed to be 160 ft3. Concrete for the base costs $3 per ft2. The materials for each of the four sides cost $2 per ft2. The flat roof will cost $7 per ft2.

OpenStudy (anonymous):

nevermind i figured it out c= 10x^2+8xh 160= (x^2)h x=4, h= 10 The minimal cost to build the shed is $480